# Key Words.: space vehicles: instruments – surveys – globular clusters: general – Galaxies: star clusters: general1 Introduction

### v2

@@
Euclid preparation.
Extragalactic globular clusters (EGCs) are an abundant and powerful tracer of galaxy dynamics and formation, and their own formation and evolution is also a matter of extensive debate. The compact nature of globular clusters means that they are hard to spatially resolve and thus study outside the Local Group. In this work we have examined how well EGCs will be detectable in images from the Euclid telescope, using both simulated pre-launch images and the first early-release observations of the Fornax galaxy cluster. The Euclid Wide Survey will provide high-spatial resolution VIS imaging in the broad band as well as near-infrared photometry (, , and ). We estimate that the galaxies within 100 Mpc in the footprint of the Euclid survey host around 830 000 EGCs of which about 350 000 are within the survey’s detection limits. For about half of these EGCs, three infrared colours will be available as well. For any galaxy within 50 Mpc the brighter half of its GC luminosity function will be detectable by the Euclid Wide Survey. The detectability of EGCs is mainly driven by the residual surface brightness of their host galaxy. We find that an automated machine-learning EGC-classification method based on real Euclid data of the Fornax galaxy cluster provides an efficient method to generate high purity and high completeness GC candidate catalogues. We confirm that EGCs are spatially resolved compared to pure point sources in VIS images of Fornax. Our analysis of both simulated and first on-sky data show that Euclid will increase the number of GCs accessible with high-resolution imaging substantially compared to previous surveys, and will permit the study of GCs in the outskirts of their hosts. Euclid is unique in enabling systematic studies of EGCs in a spatially unbiased and homogeneous manner and is primed to improve our understanding of many understudied aspects of GC astrophysics.
@@
GCs can also be used as tracers of very recent accretion events. In galaxies such as Andromeda (Mackey et al. 2013, 2019) or Centaurus A (Hughes et al. 2021), GCs have been found to align with the tidal features in the halo. This is direct evidence for the accretion and subsequent deposition of GCs hosted by a smaller galaxy into their new host. Such accreted galaxies may host nuclear star clusters in their centres that are then deposited in the galaxies’ halo where they survive typical tidal forces (Bekki & Freeman 2003; Pfeffer et al. 2016). Accreted nuclear clusters have a more extended star-formation history (Kacharov et al. 2018) and some also contain central black holes that distinguish them from normal GCs (Seth et al. 2014; Voggel et al. 2019). The correlation between nucleus mass and galaxy mass (Georgiev et al. 2009) makes remnant nuclei an excellent long-lived tracer of galaxy formation and disruption. Euclids depth and coverage with allow a census of a large area of the sky for such accretion events and candidate stripped nuclei.
@@ 2 Number of expected globular clusters for galaxies in the EWS footprint
As a first step, we thus need to identify how many local galaxies out to 100 Mpc are expected to be observed in the footprint of the planned EWS. We select all galaxies from the Heraklion Extragalactic Catalogue of local-Universe galaxies (HECATE, Kovlakas et al. 2021) that fall within the EWS region. We find a total of 24 719 galaxies that are within a distance of 100 Mpc, and have a listed magnitude (see Fig. 1). We use MOCPy (Baumann et al. 2023) to filter the list of known galaxies with the complex multi-object coverage map (MOC) of the Euclid full 6-year survey area.
The distribution of the absolute magnitudes of all selected galaxies is shown in Fig. 2. The HECATE galaxy catalogue is complete for galaxies brighter than within 33 Mpc, and out to 100 Mpc the completeness limit is (see Kovlakas et al. 2021). Thus, we are complete at the bright end of the galaxy distribution to roughly the mass of the Milky Way galaxy, which is the mass range where the vast majority of the total number of GCs is expected to reside (Harris 2016).
@@ 2.2 Number of observable GCs in Euclid
To predict the total number of GCs expected in the EWS footprint, we randomly draw a total number of GCs for each galaxy in our HECATE sample, based on its magnitude and on the distribution just described. We then randomly distribute this number of GCs among the typical magnitude distribution of GC systems, i.e., a normal distribution with a mean absolute magnitude of and a standard deviation of mag (Rejkuba 2012). The width of this GC luminosity function depends on the galaxy luminosity and is wider for brighter galaxies (Villegas et al. 2010), yet for our estimate of GC numbers that difference is negligible as other uncertainties such as the galaxy incompleteness at the faint end dominate. For each galaxy we now have a representation of the absolute magnitudes of each GC they are expected to host, and we use the distance information from the HECATE catalogue to predict how they would appear in VIS images. Figure 5 displays the distribution of apparent magnitudes of all expected GCs in the EWS footprint, for one such stochastic simulation.
@@ 3.2 Globular cluster detection in simulated VIS images
The fact that these GCs get recovered by the pipeline and included in the catalogue does not necessarily imply that we can identify them as bona fide GCs right away. For this they will have to be studied more carefully for their light profile or followed up by spectroscopy. However, the recovery fractions indicate how many we can expect in the catalogue to begin with, based on which we generate a GC candidate list.
@@ 3.3 Recovery fraction as a function of local surface brightness and GC magnitude
Figure 9 shows how the GC detectability depends on the local surface brightness at the GC location and the input magnitude of the GCs. The local surface brightness at the position of each inserted GC is derived by calculating the surface brightness in an annulus between and . The error bars shown are calculated as the 95% binomial Wilson confidence intervals.
@@
We also tested whether there was a general magnitude dependence of the recovery fraction. We find that the recovery fraction does not depend directly on GC magnitude when they are brighter than . Only when the GC magnitude is close to the VIS detection threshold we observe a sharp drop-off in the detectability of a given GC; in our simulations this is the case for our faintest GC magnitude bin, at . At all GC magnitudes brighter than this threshold, the detectability is mainly driven by the location of the GCs in terms of local surface brightness, which in the sky results from a combination of distance to the host galaxy and galaxy light profile. This emphasizes the benefits one may obtain from galaxy subtraction prior to detection, a feature that is not implemented in the standard Euclid pipeline because it is not a requirement of its core science program. The majority of GCs that we will detect with the standard pipeline will be in the outskirts of the galaxies and not within their half-light radius.
@@ 4.1 Applying a Gaussian-mixture model to select GC candidates in Fornax
The above analysis shows the efficiency of the training method in a field with known GCs. For nearby galaxies and clusters this will enable us to derive a candidate catalogue with a high completeness based solely on catalogue data.
Such a predictive modelling has the advantage that it is computationally cheap and fast, thus enabling an efficient selection of GC candidates for further study. We do expect significant contamination from both foreground and background sources, but the removal of typically more than 90% of contaminants will enable us to make a very efficient follow-up while keeping a large completeness of GCs in our sample.
@@ 6 Conclusion
If we consider a magnitude limit of for GCs we still detect the majority of a GC system within 30 Mpc. At this magnitude limit GCs will also have infrared detections in the , , and bands that are crucial to study their stellar populations and to help distinguish them from fore- and background objects.
@@ 7 Outlook
In this paper we have shown the potential that Euclid data will present for advancing globular cluster science in the future.
We estimate that there are GCs that are possibly detectable in the Euclid band. This value is based on taking into account all known galaxies within 100 Mpc in the Euclid Wide Survey area, and on estimating their total number of GCs and their magnitude distribution. This number is an estimate on how many GCs should be visible in Euclid VIS images. The real number will likely be higher due to the number of intracluster GCs that were not included in the simulation as well as the faint-end incompleteness of our input galaxy catalogue. This increase might be offset by a reduction in detected GCs due to host-galaxy background contamination, blending, and other observational effects. We also expect high-resolution VIS-imaging data for hundreds of thousands of GCs and detailed infrared colours for about half of those enabling us to produce the largest homogenous GC candidate catalogue in an non-spatially biased survey area over a third of the sky. Such a catalogue will contain many newly detected GCs, especially in the distant outskirts of galaxies and in the intracluster areas and allow for a census of GC colours, positions, luminosity function and other properties as a function of host galaxy. Such a catalogue will also allow for targeted spectroscopic follow-up of GCs that makes them usable as dynamical tracers of galaxy mass profiles and the dynamical state of entire galaxy clusters.
@@
Having used the Fornax cluster as a benchmark, we confirm that individual GCs are spatially more extended than pure point sources at 20 Mpc. This knowledge shows that a catalogue of GC light profiles and structural parameters will be possible for all sufficiently bright GCs within within 20 Mpc in the future Euclid data.
Our analysis of both simulated and first real data shows the vast number of GCs that will be present in Euclid image and will be spatially resolved and have infrared colours. We expect Euclid to be revolutionary to GC science by increasing the number of GCs accessible with space-quality high-resolution imaging by an order of magnitude or more. In particular, the wide field coverage of 14 500 deg2 enables us to search for GCs in a way that is not biased towards the centres of galaxies. We especially expect a lot of new discoveries in the outskirts of galaxies and for intracluster GCs, which have been not studied with space telescopes that have a small field-of-view.
@@ Acknowledgements.
- Cuillandre et al. (2024) Cuillandre, J.-C., Bertin, E., Bolzonella, M., et al. 2024, A&A, this issue
@@
- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, this issue
- Euclid Collaboration: Jahnke et al. (2024) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2024, A&A, this issue
- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., Achúcarro, A., et al. 2024, A&A, this issue
@@
- Marleau et al. (2024) Marleau, F., Cuillandre, J.-C., Cantiello, M., et al. 2024, A&A, this issue
@@
- Saifollahi et al. (2024) Saifollahi, T., Voggel, K., Lançon, A., et al. 2024, A&A, this issue

### v3

# Euclid preparation

@@
Euclid preparation
Extragalactic globular clusters (EGCs) are an abundant and powerful tracer of galaxy dynamics and formation, and their own formation and evolution is also a matter of extensive debate. The compact nature of globular clusters means that they are hard to spatially resolve and thus study outside the Local Group. In this work we have examined how well EGCs will be detectable in images from the Euclid telescope, using both simulated pre-launch images and the first early-release observations of the Fornax galaxy cluster. The Euclid Wide Survey will provide high-spatial resolution VIS imaging in the broad band as well as near-infrared photometry (, , and ). We estimate that the 24 719 known galaxies within 100 Mpc in the footprint of the Euclid survey host around 830 000 EGCs of which about 350 000 are within the survey’s detection limits. For about half of these EGCs, three infrared colours will be available as well. For any galaxy within 50 Mpc the brighter half of its GC luminosity function will be detectable by the Euclid Wide Survey. The detectability of EGCs is mainly driven by the residual surface brightness of their host galaxy. We find that an automated machine-learning EGC-classification method based on real Euclid data of the Fornax galaxy cluster provides an efficient method to generate high purity and high completeness GC candidate catalogues. We confirm that EGCs are spatially resolved compared to pure point sources in VIS images of Fornax. Our analysis of both simulated and first on-sky data show that Euclid will increase the number of GCs accessible with high-resolution imaging substantially compared to previous surveys, and will permit the study of GCs in the outskirts of their hosts. Euclid is unique in enabling systematic studies of EGCs in a spatially unbiased and homogeneous manner and is primed to improve our understanding of many understudied aspects of GC astrophysics.
@@
Globular clusters can also be used as tracers of very recent accretion events. In galaxies such as Andromeda (Mackey et al. 2013, 2019) or Centaurus A (Hughes et al. 2021), GCs have been found to align with the tidal features in the halo. This is direct evidence for the accretion and subsequent deposition of GCs hosted by a smaller galaxy into their new host. Such accreted galaxies may host nuclear star clusters in their centres that are then deposited in the galaxies’ halo where they survive typical tidal forces (Bekki & Freeman 2003; Pfeffer et al. 2016). Accreted nuclear clusters have a more extended star-formation history (Kacharov et al. 2018) and some also contain central black holes that distinguish them from normal GCs (Seth et al. 2014; Voggel et al. 2019). The correlation between nucleus mass and galaxy mass (Georgiev et al. 2009) makes remnant nuclei an excellent long-lived tracer of galaxy formation and disruption. Euclids depth and coverage with allow a census of a large area of the sky for such accretion events and candidate stripped nuclei.
@@ 2 Number of expected globular clusters for galaxies in the EWS footprint
As a first step, we thus need to identify how many local galaxies out to 100 Mpc are expected to be observed in the footprint of the planned EWS. We select all galaxies from the Heraklion Extragalactic Catalogue of local-Universe galaxies (HECATE, Kovlakas et al. 2021) that fall within the EWS region. We find a total of 24 719 galaxies that have a distance measurement and are within a distance of 100 Mpc, and have a listed magnitude (see Fig. 1). We can only use galaxies with a distance measurement because their absolute magnitude is required to make a prediction of how many GCs a galaxy is expected to host. GCs can be confirmed much easier in galaxies with a known distance, so this selection assures we actually predict how many GCs are detectable in the confirmed local Universe galaxies in the EWS.
We use MOCPy (Baumann et al. 2023) to filter the list of known galaxies with the complex multi-object coverage map (MOC) of the Euclid full 6-year survey area. The distribution of the absolute magnitudes of all selected galaxies is shown in Fig. 2. The HECATE galaxy catalogue is complete for galaxies brighter than within 33 Mpc, and out to 100 Mpc the completeness limit is (see Kovlakas et al. 2021). Thus, we are complete at the bright end of the galaxy distribution to roughly the mass of the Milky Way galaxy, which is the mass range where the vast majority of the total number of GCs is expected to reside (Harris 2016).
@@ 2.2 Number of observable GCs in Euclid
To predict the total number of GCs expected in the EWS footprint, we randomly draw a total number of GCs for each galaxy in our HECATE sample, based on its magnitude and on the distribution just described. The drawn Nr of GCs is then rounded to the nearest integer. In the case of the faintest galaxies this expected average GC number can be very small () and including the log-normal scatter in GC system size, it can predict GCs and thus the rounded number will be zero. The existence of dwarf galaxies that host zero GCs is naturally included in this model. The fraction of low-mass galaxies hosting no GCs is not understood observationally at this point. We do not have the statistics to deduce how common they are as we lack deep imaging needed to cover the whole GC luminosity function in a large set of dwarf galaxies at the moment. We note that with Euclid we expect progress in understanding the GC luminosity function (or the absence) in dwarfs and first observational results have been shown in (Saifollahi et al. 2024). As the number of GCs in these very faint galaxies is negligible compared to the total expected number the effect of this approximation is very small and within the uncertainties of the final forecast number.
We then randomly distribute this number of GCs among the typical magnitude distribution of GC systems, i.e., a normal distribution with a mean absolute magnitude of and a standard deviation of mag (Rejkuba 2012). The width of this GC luminosity function depends on the galaxy luminosity and is wider for brighter galaxies (Villegas et al. 2010), yet for our estimate of GC numbers that difference is negligible as other uncertainties such as the galaxy incompleteness at the faint end dominate. For each galaxy we now have a representation of the absolute magnitudes of each GC they are expected to host, and we use the distance information from the HECATE catalogue to predict how they would appear in VIS images. Figure 5 displays the distribution of apparent magnitudes of all expected GCs in the EWS footprint, for one such stochastic simulation.
@@ 3.2 Globular cluster detection in simulated VIS images
The fact that these GCs get recovered by the pipeline and included in the catalogue does not necessarily imply that we can identify them as bona fide GCs right away. For this they will have to be studied more carefully for their light profile or followed up by spectroscopy. However, the recovery fractions indicate how many we can expect in the catalogue to begin with, based on which we generate a GC candidate list. Those GCs that are not directly recovered by the Euclid photometric pipeline will still exist in the images if they are brighter than the detection threshold. The large spatial coverage of the Euclid survey means that detecting them in a systematic manner will be challenging because it requires treating each image separately and for example removing the light of the surrounding galaxy.
@@ 3.3 Recovery fraction as a function of local surface brightness and GC magnitude
The left panel of Fig. 9 shows how the GC detectability depends on the local surface brightness at the GC location and the input magnitude of the GCs. The local surface brightness at the position of each inserted GC is derived by calculating the surface brightness in an annulus between and . The error bars shown are calculated as the 95% binomial Wilson confidence intervals.
@@
We also tested whether there was a general magnitude dependence of the recovery fraction. The overall detectability is shown in the right panel of Fig. 9, where we find that the recovery fraction does not depend directly on GC magnitude when they are brighter than . Only when the GC magnitude is close to the VIS detection threshold we observe a sharp drop-off in the detectability of a given GC; in our simulations this is the case for our faintest GC magnitude bin, at . At all GC magnitudes brighter than this threshold, the detectability is mainly driven by the location of the GCs in terms of local surface brightness, which in the sky results from a combination of distance to the host galaxy and galaxy light profile. This emphasizes the benefits one may obtain from galaxy subtraction prior to detection, a feature that is not implemented in the standard Euclid pipeline because it is not a requirement of its core science program. The majority of GCs that we will detect with the standard pipeline will be in the outskirts of the galaxies and not within their half-light radius.
@@ 4.1 Applying a Gaussian-mixture model to select GC candidates in Fornax
The above analysis shows the efficiency of the training method in a field with known GCs. For nearby galaxies and clusters this will enable us to derive a candidate catalogue with a high completeness based solely on catalogue data. Such a predictive modelling has the advantage that it is computationally cheap and fast, thus enabling an efficient selection of GC candidates for further study. We do expect significant contamination from both foreground and background sources, but the removal of typically more than 90% of contaminants will enable us to make a very efficient follow-up while keeping a large completeness of GCs in our sample.
@@ 6 Conclusion
GCs will also be detected in the infrared , , and bands down to a magnitude limit of . Infrared colors are crucial to study a GCs stellar populations and to help distinguish them from fore- and background objects. This magnitude limit implies that the brightest half of a GC system of a given galaxy is detectable in the infrared out to 30 Mpc of distance.
@@ 7 Outlook
In this paper we have shown the potential that Euclid data will present for advancing globular cluster science in the future. We estimate that there are GCs that are possibly detectable in the Euclid band. This value is based on taking into account all known galaxies within 100 Mpc in the Euclid Wide Survey area, and on estimating their total number of GCs and their magnitude distribution. This number is an estimate on how many GCs should be visible in Euclid VIS images. The real number will likely be higher due to the number of intracluster GCs that were not included in the simulation as well as the faint-end incompleteness of our input galaxy catalogue. This increase might be offset by a reduction in detected GCs due to host-galaxy background contamination, blending, and other observational effects.
We also expect high-resolution VIS-imaging data for hundreds of thousands of GCs and detailed infrared colours for about half of those enabling us to produce the largest homogenous GC candidate catalogue in an non-spatially biased survey area over a third of the sky. Such a catalogue will contain many newly detected GCs, especially in the distant outskirts of galaxies and in the intracluster areas and allow for a census of GC colours, positions, luminosity function and other properties as a function of host galaxy. Such a catalogue will also allow for targeted spectroscopic follow-up of GCs that makes them usable as dynamical tracers of galaxy mass profiles and the dynamical state of entire galaxy clusters.
@@
Having used the Fornax cluster as a benchmark, we confirm that individual GCs are spatially more extended than pure point sources at 20 Mpc. This knowledge shows that a catalogue of GC light profiles and structural parameters will be possible for all sufficiently bright GCs within within 20 Mpc in the future Euclid data. Our analysis of both simulated and first real data shows the vast number of GCs that will be present in Euclid image and will be spatially resolved and have infrared colours. We expect Euclid to be revolutionary to GC science by increasing the number of GCs accessible with space-quality high-resolution imaging by an order of magnitude or more. In particular, the wide field coverage of 14 500 deg2 enables us to search for GCs in a way that is not biased towards the centres of galaxies. We especially expect a lot of new discoveries in the outskirts of galaxies and for intracluster GCs, which have been not studied with space telescopes that have a small field-of-view.
@@ Acknowledgements.
- Cuillandre et al. (2024) Cuillandre, J.-C., Bertin, E., Bolzonella, M., et al. 2024, A&A, submitted, arXiv:2405.13496
@@
- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, accepted, arXiv:2405.13492
- Euclid Collaboration: Jahnke et al. (2024) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2024, A&A, accepted, arXiv:2405.13493
- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2024, A&A, accepted, arXiv:2405.13491
@@
- Marleau et al. (2024) Marleau, F., Cuillandre, J.-C., Cantiello, M., et al. 2024, A&A, accepted, arXiv:2405.13502
@@
- Saifollahi et al. (2024) Saifollahi, T., Voggel, K., Lançon, A., et al. 2024, A&A, accepted, arXiv:2405.13500